# Generalizable Implicit Motion Modeling for Video Frame Interpolation

### v3

@@ 4.2 Ablation Studies
Latent refinement. To validate the efficacy of latent refinement in GIMM, we conducted experiments excluding the latent refiner (Non-Refiner). This led to notable declines of 0.53 dB and 0.43 dB in PSNRs across both benchmarks. Thus, refining the motion latent proves essential for accurate motion modeling.
Spatial coordinates. We use 3D spatiotemporal coordinates in the implicit modeling of GIMM. To assess the necessity of spatial coordinates, we replace 3D coordinates with temporal coordinates (T-coord only). This removal causes a 0.16 dB drop of PSNR on Vimeo-Triplet-Flow and a 0.06 increase of EPE on Vimeo-Septuplet-Flow. This emphasizes the crucial role of spatial coordinates in implicit modeling.
@@
RIFE [19] 31.43† 30.58† 36.33 31.87 27.21
IFRNet [25] 31.53† 30.46† 34.88 31.15 26.32
M2M [17] 32.13† 30.88† 36.56 31.92 27.14
AMT [28] 28.88 28.17 34.49 31.03 26.44
UPR-Net [22] 31.16 30.50 36.78 31.96 27.14
EMA-VFI [52] 32.53 31.21 36.65 32.28 27.72
CURE [42] 30.24 - 36.09 31.32 26.61
GIMM-VFI-R 32.71 31.88 37.02 32.62 27.99
GIMM-VFI-F 32.91 31.97 37.03 32.56 28.01
@@ 4.3 Arbitrary-timestep Video Frame Interpolation
Evaluation settings. We calculate PSNR on both SNU-FILM-arb and X4K-1000FPS (XTest) [43]. XTest comprises 15 clips from 15 videos for 8X interpolation. Following the established protocol from prior studies [17, 52], we assess interpolation quality at both 2K and 4K resolutions. For fair comparisons, we disable test-time augmentations during evaluations.
@@
Results. We present the quantitative results of GIMM-VFI on arbitrary-timestep interpolation benchmarks in Table 3. Our proposed method achieves high-quality continuous interpolation across various timesteps (i.e., 4X, 8X, 16X). We observe that our RAFT-based [47] method GIMM-VFI-R achieves significant improvements of 0.18 dB on XTest-2K, 0.67 dB on XTest-4K, and approximately 0.30 dB on each subset of the SNU-FILM-arb, in comparison with the previous state-of-the-art method EMA-VFI [52]. This demonstrates that our GIMM can offer effective motion modeling for the arbitrary-timestep VFI task.
@@
5 Conclusion
We propose a novel motion modeling method, GIMM, which performs continuous motion modeling for video frame interpolation. Our proposed GIMM is the first attempt to learn generalizable implicit neural representations for continuous motion modeling. The method can properly predict arbitrary-timestep optical flows within two adjacent video frames at any resolution, and be easily integrated with existing flow-based VFI approaches(e.g., AMT [28]) without further modifications. Extensive experiments on the VFI benchmarks show that our GIMM can effectively perform generalizable motion modeling across videos.

### v4

@@ 4.3 Ablation Study on GIMM
Motion encoder. We leverage the motion encoder in our GIMM to extract features from flows, alleviating the negative impacts brought by the possible bias and noises in the estimated flows. To justify this design, we remove the motion encoder for a direct comparison (Non-ME). As a result, we observe increases in error on both benchmarks, e.g., the EPE on VSF rises 0.17. Therefore, the motion encoder benefits the motion modeling in GIMM.
Latent refinement. To validate the efficacy of latent refinement in GIMM, we conducted experiments excluding the latent refiner (Non-Refiner). This leads to notable declines of 0.53 dB and 0.43 dB in PSNRs across both benchmarks. Thus, refining the motion latent proves essential for accurate motion modeling.
Spatial coordinates. Since the modeled motion consists of spatiotemporal changes, we use 3D spatiotemporal coordinates in the implicit modeling of GIMM. To assess the necessity of spatial coordinates, we replace 3D coordinates with temporal coordinates (T-coord only). This removal causes a 0.16 dB drop of PSNR on Vimeo-Triplet-Flow and a 0.06 increase of EPE on Vimeo-Septuplet-Flow. This emphasizes the crucial role of spatial coordinates in implicit modeling.
@@
RIFE [20] 31.43† / 0.126 / 11.99 30.58† / 0.152 / 13.52 36.33 / 0.038 / 6.65 31.87 / 0.072 / 11.99 27.21 / 0.134 / 19.82
M2M [18] 32.13† / 0.098 / 9.25 30.88† / 0.158 / 8.67 36.56 / 0.036 / 5.98 31.92 / 0.061 / 10.13 27.14 / 0.112 / 17.37
IFRNet [26] 31.53† / 0.108 / 23.93 30.46† / 0.164 / 23.75 34.88 / 0.046 / 9.92 31.15 / 0.066 / 11.65 26.32 / 0.115 / 16.91
AMT [29] 28.88 / 0.153 / 13.92 28.17 / 0.187 / 13.97 34.49 / 0.072 / 9.25 31.03 / 0.089 / 10.34 26.44 / 0.136 / 14.72
UPR-Net [23] 31.16 / 0.104 / 10.75 30.50 / 0.154 / 9.45 36.78 / 0.033 / 6.09 31.96 / 0.064 / 9.93 27.14 / 0.111 / 16.76
EMA-VFI [55] 32.53 / 0.097 / 7.21 31.21 / 0.156 / 8.61 36.65 / 0.041 / 7.07 32.28 / 0.074 / 12.17 27.72 / 0.130 / 19.58
CURE [45] 30.24 / 0.111 / 26.42 - 36.09 / 0.035 / 6.98 31.32 / 0.063 / 12.72 26.61 / 0.114 / 22.62
GIMM-VFI-R 32.71 / 0.113 / 6.52 31.88 / 0.149 / 6.49 37.02 / 0.033 / 5.89 32.62 / 0.060 / 9.59 27.99 / 0.110 / 16.45
GIMM-VFI-F 32.91 / 0.103 / 6.74 31.97 / 0.142 / 6.58 37.03 / 0.031 / 5.86 32.56 / 0.059 / 9.95 28.01 / 0.109 / 16.79
@@ 4.4 Arbitrary-timestep Video Frame Interpolation
Evaluation settings. We calculate PSNR and perceptual indicators, e.g., LPIPS [56], FID [16], on both SNU-FILM-arb and X4K-1000FPS (XTest) [46]. XTest comprises 15 clips from 15 videos for 8X interpolation. Following the established protocol from prior studies [18, 55], we assess interpolation quality at both 2K and 4K resolutions. For fair comparisons, we disable test-time augmentations during evaluations. We also provide additional evaluations of our method for perception-oriented interpolation in the supplementary materials.
@@
Results. We present the quantitative results of GIMM-VFI on arbitrary-timestep interpolation benchmarks in Table 4. Our proposed method achieves high-quality continuous interpolation across various timesteps (i.e., 4X, 8X, 16X). In terms of PSNR, we observe that our RAFT-based [50] method GIMM-VFI-R achieves significant improvements of 0.18 dB on XTest-2K, 0.67 dB on XTest-4K, and approximately 0.30 dB on each subset of the SNU-FILM-arb, in comparison with the previous state-of-the-art method EMA-VFI [55]. For perceptual metrics, i.e., LPIPS and FID, our proposed GIMM-VFI still outperforms previous methods on most benchmarks, delivering competitive performance. These observations demonstrate that our GIMM can offer effective motion modeling for the arbitrary-timestep VFI task.
@@ 5 Limitation
6 Conclusion
We propose a novel motion modeling method, GIMM, which performs continuous motion modeling for video frame interpolation. Our proposed GIMM is the first attempt to learn generalizable implicit neural representations for continuous motion modeling. The method can properly predict arbitrary-timestep optical flows within two adjacent video frames at any resolution, and be easily integrated with existing flow-based VFI approaches (e.g., AMT [29]) without further modifications by supplying accurately modeled motion. Extensive experiments on the VFI benchmarks show that our GIMM can effectively perform generalizable motion modeling across videos.
Acknowledgement. This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contributions from the industry partner(s).